# Enhancing Physics-Informed Neural Networks Through Feature Engineering

### v2

@@ 4 Related work
Architectural Modifications. One way to improve the training procedure for PINNs is to modify the network architecture, improving the optimization landscape relative to the basic PINN, making training easier. Examples of this approach include the adapative activation functions of (Jagtap et al., 2020) (A-PINNs), and specialized architectures designed to mitigate spectral bias (Li et al., 2020a).
Loss Re-weighting. Another popular technique is loss reweighting. For certain PDEs, the residual loss tends to dominate the boundary loss in that the optimizer focuses too much on minimizing the residual loss, leading to a solution that fails to satisfy the boundary conditions. To address this, techniques like W-PINNs (Wang et al., 2021a) balance loss components through heuristic or learned weights to down-weight the residual loss and better fit the boundary loss.
Optimizer Design. Another popular approach to PINN training is to develop more sophisticated optimizers that are more robust to ill-conditioning. Several notable proposals in this area are the natural gradients method of (Müller & Zeinhofer, 2023), MultiAdam (Yao et al., 2023), and NysNewton-CG (Rathore et al., 2024). These methods target ill-conditioning directly and use curvature information from the loss or the model to precondition the gradient. This leads to an improved optimization landscape locally, enabling the optimizer to take better steps and progress faster. In contrast, SAFE-NET enjoys an implicit preconditioning effect that is global—by incorporating a trainable feature layer, SAFE-NET changes the PINN objective, globally changing the optimization landscape. The results in Section 5.3 show SAFE-NET enjoys a significantly better-conditioned optimization landscape. Thus, SAFE-NET can be further combined with more sophisticated optimization schemes to obtain further improvements.
@@ 5.1 Baselines
We test against PINN models with feature engineering (RBF, RBF-P, RFF, and RBA-PINN) and without (PINN, FLS-PINN, and W-PINN). Details on each baseline appear in Appendix A.2.
@@
We experiment on eight PDE tasks as provided in Table 2. Depending on availability, the datasets for these tasks are either from PDE benchmarks such as PDEBench (Takamoto et al., 2024) and PINNacle (Hao et al., 2023) or implemented directly if unavailable online. More details on each PDE and its source are provided in Appendix B.
Table 2: Overview of the tested PDEs.
@@ 5.2 Experiments
Following prior work (such as (Rathore et al., 2024) and (Zeng et al., 2024b)), we use a two-stage Adam + L-BFGS optimizer for the first set of experiments. Recent loss-landscape studies on PINNs show that first-order methods such as Adam make rapid progress away from saddle points but eventually slow down in the ill-conditioned valleys induced by differential operators, whereas quasi-Newton methods like L-BFGS improve local conditioning yet are attracted to saddles and often stall when started from scratch. Combining them inherits the best of both worlds: an Adam “warm-up” steers the parameters clear of saddles; switching to L-BFGS then preconditions the Hessian and yields much faster local convergence . Nevertheless, pure L-BFGS can terminate early because its strong-Wolfe line search sometimes returns a step size of zero, especially on stiff PDEs . Our schedule therefore runs Adam first until progress slows down significantly and switches to L-BFGS until the maximum number of iterations is reached. A more detailed theoretical justification of the Adam + L-BFGS schedule appears in (Rathore et al., 2024).
@@
Table 3: PDE benchmark results comparing several PINN architectures in relative $L^{2}$ error. The best results are shown in Bold. However, numerically close results (i.e., within 30% of the best result, including the best result in Bold itself) are also highlighted in Blue. This accounts for potential variance in implementation and hyperparameter sensitivity. Cells with (*) indicate L-BFGS divergence in our experiments. PDE abbreviations: W=Wave, R=Reaction, D=Diffusion, H=Heat, C=Convection, A-C=Allen-Cahn, B=Burgers, N-S=Navier-Stokes.
@@
W $6.62\text{e}{-2}$ $4.37\text{e}{-2}$ $8.79\text{e}{-3}$ $7.39\text{e}{-3}$
R $4.98\text{e}{-2}$ $1.27\text{e}{-1}$ $4.52\text{e}{-2}$ $1.20\text{e}{-2}$
D $9.65\text{e}{-3}$ $5.89\text{e}{-2}$ $1.43\text{e}{-3}$ $2.56\text{e}{-3}$
H $6.34\text{e}{-3}$ $6.98\text{e}{-3}$ $9.98\text{e}{-3}$ $8.56\text{e}{-3}$
C $1.96\text{e}{-2}$ $4.12\text{e}{-2}$ $9.77\text{e}{-3}$
A-C $4.98\text{e}{-1}$ $1.46\text{e}{-0}$ $1.29\text{e}{-1}$ $1.03\text{e}{-2}$
B $1.01\text{e}{-2}$ $6.16\text{e}{-2}$*}
N-S $7.49\text{e}{-1}$ $7.91\text{e}{-1}$ $6.76\text{e}{-1}$
@@
W $1.57\text{e}{-3}$ $2.48\text{e}{-2}$ $2.13\text{e}{-2}$ $\boldsymbol{1.21\text{e}{-3}}$
R $3.31\text{e}{-2}$ $1.98\text{e}{-2}$ $1.37\text{e}{-2}$ $\boldsymbol{9.93\text{e}{-3}}$
D $1.29\text{e}{-4}$ ${4.02\text{e}{-4}}$ $\boldsymbol{9.91\text{e}{-5}}$ ${1.21\text{e}{-4}}$
H $7.35\text{e}{-4}$ $3.65\text{e}{-3}$ $7.21\text{e}{-4}$ $\boldsymbol{5.31\text{e}{-4}}$
C $8.32\text{e}{-3}$ $7.72\text{e}{-2}$ $7.02\text{e}{-3}$ $\boldsymbol{4.37\text{e}{-3}}$
A-C $\boldsymbol{8.21\text{e}{-5}}$ $1.06\text{e}{-4}$ ${1.97\text{e}{-3}}$
B $4.17\text{e}{-4}$ ${3.95\text{e}{-3}}$ $\boldsymbol{3.47\text{e}{-4}}$ $2.67\text{e}{-3}$
N-S ${4.86\text{e}{-1}}$ $\boldsymbol{2.56\text{e}{-1}}$ ${2.98\text{e}{-1}}$ $5.26\text{e}{-1}$
Even after warming up with Adam, L-BFGS can still cause divergence or instability on some PDEs and baseline methods. Corresponding entries are denoted by “*’ in the result table. There could be several reasons why this is possible. For instance, RFF uses fixed random frequencies drawn from a Gaussian distribution, which can lead to very poor conditioning depending on the random initialization. The mismatch between random frequencies and the true solution frequencies can create ill-conditioned optimization landscapes that cause L-BFGS to diverge. Alternatively, in the case of W-PINN and Navier-Stokes, W-PINN’s adaptive weighting scheme computes weights based on NTK matrix traces (see (Wang et al., 2022b) for the details or the computation). Due to the complex flow physics of Navier-Stokes, the NTK matrices could become ill-conditioned near boundary layers or rapid changes in the adaptive weights can destabilize L-BFGS.
SAFE-NET, on the other hand, is compatible with quasi-Newton methods such as L-BFGS: it can safely be optimized by them. Table 3 demonstrates that for shock-free PDEs (i.e., not Burgers or Navier-Stokes, and Allen-Cahn to some extent), SAFE-NET outperforms competing methods in L2RE by in 4/5 tasks, following closely behind RBF-P for the diffusion task. (See Figure 3 for visual demonstration). Allen-Cahn is “less nonlinear” than Burgers or Navier-Stokes, but can also develop steep gradients and interfaces, especially in phase separation dynamics. In practice, these sharp transition layers are not true shocks in the hyperbolic PDE sense, but numerically they still present challenges similar to shocks — including Gibbs phenomena for Fourier features in particular. Despite that, SAFE-NET manages to come third best after RBA-PINN and RBF-P for both Allen-Cahn and Burgers tasks. This outcome is not surprising as SAFE-NET is not designed for PDEs with shocks. Particularly, (Zeng et al., 2024b) suggests that the RBF kernel handles the discontinuity at $x=0$ in the Burgers equation more effectively.
@@
[…]
@@
[…]
Figure 8: Average percentage of improvement or deterioration (in L2RE) for each method after switching from Optimization Schedule (1) to Optimization Schedule (2). Figure 12 in the appendix shows precentages for every method on every PDE, demonstrating that SAFE-NET is the only method that produces results in consistent improvement on every PDE after switching to (Adam + L-BFGS)2.
@@ 5.3 Feature Engineering & Spectral Density Analysis
SAFE-NET significantly improves the conditioning of all tested PDEs. We analyze conditioning both at initialization (3k epochs) and after training for a while (100k iterations) using SAFE-NET and PINN with Adam as the optimizer with an initial learning rate of 0.001 and an exponential decay rate of 0.9 every 2k steps (Experimental setup of Figure 1)
@@
We see that SAFE-NET shifts most eigenvalues values close to 1 at initialization (Figure 10 top). Post-training spectral density plots (Figure 10 bottom) reveal dramatic conditioning improvements: the top eigenvalues for the wave problem decrease by a factor of $10^{4}$, while those for the heat and Burgers problems decrease by $10^{2}$ (Figures 24 and 25. Further experiments (Appendix C) show that SAFE-NET reduces both the number and density of large eigenvalues across all problems and all baseline methods (see Table 20)
@@
This improvement in conditioning is a key factor explaining why SAFE-NET features are easier to optimize. Figure 11 compares SAFE-NET and PINNs performance on different loss components for the diffusion and Allen-Cahn PDEs (from the same experiment as Figure 5) as an example. (Rathore et al., 2024) argues the residual loss is the main cause of the ill-conditioned loss landscape of PINNs. SAFE-NET significantly improves the conditioning of each loss component, including the residual loss.
@@ 6 Implicit preconditioning in SAFE-NET
Empirical results in Section 5.3 show SAFE-NET improves the conditioning of each problem both at the beginning and end of training. To develop a better intuition for why this is the case, we consider a simple didactic setting similar to Wang et al. (2020), where the network is given by
@@
$\Theta_{f}(x^{\prime},x)=\nabla_{\theta}f(x)^{\top}\nabla f_{\theta}(x).$
@@
where the expectation is taken over the weights at initialization (Jacot et al., 2018; Liu et al., 2020). The neural tangent kernel matrix $\Theta^{\infty}_{f}$ is defined analogously to the tangent kernel matrix.
In the limit $p\rightarrow\infty$, neural net training with $f_{\theta}$ is equivalent to kernel regression with the NTK matrix $\Theta^{\infty}_{f}$. As training is reduced to a kernel regression problem, the convergence speed of gradient-based optimizers is controlled by the conditioning of the NTK matrix $\Theta^{\infty}_{f}$ (Jacot et al., 2018; Liu et al., 2022). Thus, a better-conditioned NTK yields a better optimization landscape and faster convergence.
[…]
@@
$\phi(x,t)=\sum_{i=1}^{N}\sum_{j=1}^{4}c_{j}^{(i)}\phi_{j}^{(i)}(x,t),$
where $\{\phi_{j}^{(i)}\}_{j=1}^{4}$ are defined in equations (5)-(8) and $i=1,...,N$ are the number of each type of product features. The kernel is a sum of separable functions as
$\Theta_{u}((x,t),(x^{\prime},t^{\prime}))=\sum_{i=1}^{N}\sum_{j=1}^{4}\big{(}c_{j}^{(i)}\big{)}^{2}\phi_{j}^{(i)}(x,t)\phi_{j}^{(i)}(x^{\prime},t^{\prime}).$ (12)
@@
[…]
Substituting $\Theta_{n}((x,t),(x^{\prime},t^{\prime}))$ from equation (12), we get
$\sum_{i=1}^{N}\sum_{j=1}^{4}\big{(}c_{j}^{(i)}\big{)}^{2}\phi_{j}^{(i)}(x,t)\int\phi_{j}^{(i)}(x^{\prime},t^{\prime})\psi(x^{\prime},t^{\prime})=\beta\psi(x,t).$ (13)
[…]
$\psi(x,t)=\frac{1}{\beta}\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(i)}\big{(}c_{j}^{(i)}\big{)}^{2}\phi_{j}^{(i)}(x,t).$ (14)
@@
$\frac{1}{\beta}\sum_{i,\ell}^{N}\sum_{j,m}^{4}\big{(}c_{j}^{(i)}\big{)}^{2}\phi_{j}^{(i)}\alpha_{m}^{(\ell)}\big{(}c_{m}^{(\ell)}\big{)}^{2}\int\phi_{j}^{(i)}(x^{\prime},t^{\prime})\phi_{m}^{(\ell)}(x^{\prime},t^{\prime})$
$=\beta\big{(}\frac{1}{\beta}\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(i)}\big{(}c_{j}^{(i)}\big{)}^{2}\phi_{j}^{(i)}\big{)}$
Now, $\{\phi_{j}^{(i)}\}$ forms an orthonormal basis, as each $\phi$ is chosen from an orthogonal Fourier basis and has unit norm. Thus,
$\frac{1}{\beta}\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(i)}\big{(}c_{j}^{(i)}\big{)}^{4}\phi_{j}^{(i)}=\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(i)}\big{(}c_{j}^{(i)}\big{)}^{2}\phi_{j}^{(i)}.$
From this display, we conclude that for each pair $(i,j)$, $\alpha_{j}^{(i)}\left(c_{j}^{(i)}\right)^{2}=0\quad\text{or}\quad\beta=\left(c_{j}^{(i)}\right)^{2}.$ Thus, the eigenvalues of $T_{\Theta_{u}}$ are 00 or equal $(c_{j}^{(i)})^{2}$ with corresponding eigenfunction $\psi(x,t)=\alpha_{j}^{(i)}\phi_{j}^{(i)}(x,t)$. As the non-zero eigenvalues correspond to the directions relevant to learning, we focus on them. Recall the $\{c_{j}^{(i)}\}$ are the trainable amplitudes and are set to $1$ at initialization. Hence, at initialization, we expect the condition number of the matrix $\Theta_{u}$ to be approximately 1. If $u$ were in the infinite width limit, this would imply $\Theta^{\infty}_{u}$ is well-conditioned and fast convergence of gradient-based optimizers. Thus, this idealized example shows the features selected by SAFE-NET can lead to better conditioning and faster convergence. In practice, finite network width and perturbations from domain knowledge features and normalization are expected to introduce some spreading in the eigenvalue distribution. However, experiments show that SAFE-NET maintains a denser eigenvalue distribution around the theoretical prediction from the example (see Figure 10(a)), preserving the core conditioning benefits anticipated by the idealized example and its insights. Early on in training, Figure 10 shows the eigenvalue distribution is relatively uniform. During training, SAFE-NET gradually adjusts the $c_{j}^{(i)}$’s, balancing different frequencies rather than letting any single Fourier mode dominate. As a result, the eigenvalue distribution shifts outwards relatively slowly. Thus, even at the end of the training, Figure 10 shows the eigenvalue distribution has not changed much from its initialization, so the landscape remains well-conditioned.

### v3

@@ 4 Related work
Architectural Modifications. One way to improve the training procedure for PINNs is to modify the network architecture, improving the optimization landscape relative to the basic PINN, making training easier. Examples of this approach include the adapative activation functions of jagtap2020adaptive (A-PINNs), and specialized architectures designed to mitigate spectral bias (li2020d3m).
Loss Re-weighting. Another popular technique is loss reweighting. For certain PDEs, the residual loss tends to dominate the boundary loss in that the optimizer focuses too much on minimizing the residual loss, leading to a solution that fails to satisfy the boundary conditions. To address this, techniques like W-PINN wang2022and balance loss components through heuristic or learned weights to down-weight the residual loss and better fit the boundary loss.
Optimizer Design. Another popular approach to PINN training is to develop more sophisticated optimizers that are more robust to ill-conditioning. Several notable proposals in this area are the natural gradients method of muller2023achieving, MultiAdam yao2023multiadam, and NysNewton-CG rathore2024challenges. These methods target ill-conditioning directly and use curvature information from the loss or the model to precondition the gradient. This leads to an improved optimization landscape locally, enabling the optimizer to take better steps and progress faster. In contrast, SAFE-NET enjoys an implicit preconditioning effect that is global—by incorporating a trainable feature layer, SAFE-NET changes the PINN objective, globally changing the optimization landscape. The results in Section˜5.3 show SAFE-NET enjoys a significantly better-conditioned optimization landscape. Thus, SAFE-NET can be further combined with more sophisticated optimization schemes to obtain further improvements.
@@ 5.1 Baselines
We compare two groups of PINN baselines: (i) without feature engineering (PINN, FLS-PINN, W-PINN) and (ii) with feature engineering (RFF, RBF, RBF-P, RBA-PINN). Details on each baseline could be found in Appendix A.2.
@@
We experiment on eight PDE benchmarks in this section as provided in Table˜2. Depending on availability, the datasets for these tasks are either from PDE benchmarks such as PDEBench takamoto2024pdebenchextensivebenchmarkscientific and PINNacle hao2023pinnacle or implemented directly if unavailable online. More details on each PDE and its source are provided in Appendix B. An additional experiment on a synthesized non-homogeneous boundary heat task is included in Appendix C.3.
Table 2: Overview of the tested PDEs. Following the conventions of PINNs literature, the dimensions provided in this table are spatial dimensions. An extra dimension should be added for time-dependent PDEs to reach the correct number of variables.
@@ 5.2 Experiments
Following prior work (such as rathore2024challenges and zeng2024rbfpinnnonfourierpositionalembedding), we use a two-stage Adam + L-BFGS optimizer for the first set of experiments. Recent loss-landscape studies on PINNs show that first-order methods such as Adam make rapid progress away from saddle points but eventually slow down in the ill-conditioned valleys induced by differential operators, whereas quasi-Newton methods like L-BFGS improve local conditioning yet are attracted to saddles and often stall when started from scratch. Combining them inherits the best of both worlds: an Adam “warm-up” steers the parameters clear of saddles; switching to L-BFGS then preconditions the Hessian and yields much faster local convergence . Nevertheless, pure L-BFGS can terminate early because its strong-Wolfe line search sometimes returns a step size of zero, especially on stiff PDEs . Our schedule therefore runs Adam first until progress slows down significantly and switches to L-BFGS until the maximum number of iterations is reached. A more detailed theoretical justification of the Adam + L-BFGS schedule appears in rathore2024challenges.
@@
Table 3: PDE benchmark results comparing several PINN architectures in relative $L^{2}$ error. The best results are shown in Bold. However, numerically close results (i.e., within 30% of the best result, including the best result in Bold itself) are also highlighted in Blue. This accounts for potential variance in implementation and hyperparameter sensitivity. Cells with (*) indicate L-BFGS divergence in our experiments.
@@
Wave $6.62\text{e}{-2}$ $4.37\text{e}{-2}$ $8.79\text{e}{-3}$ $7.39\text{e}{-3}$
Reaction $4.98\text{e}{-2}$ $1.27\text{e}{-1}$ $4.52\text{e}{-2}$ $1.20\text{e}{-2}$
Diffusion $9.65\text{e}{-3}$ $5.89\text{e}{-2}$ $1.43\text{e}{-3}$ $2.56\text{e}{-3}$
Heat $6.34\text{e}{-3}$ $6.98\text{e}{-3}$ $9.98\text{e}{-3}$ $8.56\text{e}{-3}$
Convection $1.96\text{e}{-2}$ $4.12\text{e}{-2}$ $9.77\text{e}{-3}$ *
Allen-Cahn $4.98\text{e}{-1}$ $1.46\text{e}{-0}$ $1.29\text{e}{-1}$ $1.03\text{e}{-2}$
Burgers $1.01\text{e}{-2}$ $6.16\text{e}{-2}$ * {*}
Navier-Stokes $7.49\text{e}{-1}$ $7.91\text{e}{-1}$ * $6.76\text{e}{-1}$
@@
[…]
Even after warming up with Adam, L-BFGS can still cause divergence or instability on some PDEs and baseline methods. Corresponding entries are denoted by “*’ in the result table. There could be several reasons why this is possible. For instance, RFF uses fixed random frequencies drawn from a Gaussian distribution, which can lead to very poor conditioning depending on the random initialization. The mismatch between random frequencies and the true solution frequencies can create ill-conditioned optimization landscapes that cause L-BFGS to diverge. Alternatively, in the case of W-PINN and Navier-Stokes, W-PINN’s adaptive weighting scheme computes weights based on NTK matrix traces (see wang2022and for the details or the computation). Due to the complex flow physics of Navier-Stokes, the NTK matrices could become ill-conditioned near boundary layers or rapid changes in the adaptive weights can destabilize L-BFGS.
SAFE-NET, on the other hand, is compatible with quasi-Newton methods such as L-BFGS: it can safely be optimized by them. Table 3 demonstrates that for shock-free PDEs (i.e., not Burgers or Navier-Stokes, and Allen-Cahn to some extent), SAFE-NET outperforms competing methods in L2RE by in 4/5 tasks, following closely behind RBF-P for the diffusion task. (See Figure˜3 for visual demonstration). Allen-Cahn is “less nonlinear" than Burgers or Navier-Stokes, but can also develop steep gradients and interfaces, especially in phase separation dynamics. In practice, these sharp transition layers are not true shocks in the hyperbolic PDE sense, but numerically they still present challenges similar to shocks — including Gibbs phenomena for Fourier features in particular. Despite that, SAFE-NET manages to come third best after RBA-PINN and RBF-P for both Allen-Cahn and Burgers tasks. This outcome is not surprising as SAFE-NET is not designed for PDEs with shocks. Particularly, zeng2024rbfpinnnonfourierpositionalembedding suggests that the RBF kernel handles the discontinuity at $x=0$ in the Burgers equation more effectively.
@@
Figure 5 demonstrates the performance of SAFE-NET + $(\text{Adam}+\text{L-BFGS})^{2}$ on the diffusion and Allen-Cahn PDEs, presenting the gradient norm plots as well. Table 4 summarizes these numerical results, showing significant improvements on more challenging PDEs containing steep interface transitions, shocks, and discontinuities (Allen-Cahn, Burgers, and Navier-Stokes) along with improvements on other PDEs matching the top results of Table˜3. These results demonstrate the effectiveness of $(\text{Adam}+\text{L-BFGS})^{2}$ when paired with SAFE-NET, leading to significant improvements compared with Adam + L-BFGS, especially in the diffusion and Allen-Cahn tasks. With this improved optimizer, SAFE-NET beats all compared architectures on 5/8 of the tested PDEs as well as showing improvements for all the nonlinear PDEs with shocks; see Figure 6 for a visual demonstration.
Figure 6: Comparison of SAFE-NET + $(\text{Adam}+\text{L-BFGS})^{2}$ results for each PDE against the best baseline performance achieved using either (Adam + L-BFGS) or $(\text{Adam}+\text{L-BFGS})^{2}$. This comparison selects the optimal result for each competing method across both optimization strategies, highlights SAFE-NET’s consistent advantages on shock-free PDEs (now ranking first on 5/8 tasks) and validates the effectiveness of $(\text{Adam}+\text{L-BFGS})^{2}$ with SAFE-NET. Compare with Figure 3.
@@
Tables 5 and 6 appearing in Appendix A.3 contain complete numerical results for Optimization Schedule (1) and (2) along with other combinations tested. Notably, $(\text{Adam}+\text{L-BFGS})^{2}$ fails to improve (and often degrades) performance for the majority of the baseline methods across different PDEs, making it unsuitable to compare PINN architectures. Figure 8 shows the average percentage of improvement or deterioration (in L2RE) for each method after switching from Optimization Schedule (1) to Optimization Schedule (2). To see an example of the loss curves, Figure 7 compares the performances of the top 3 best performing methods on the diffusion PDE using Adam + L-BFGS (Optimization Schedule (1)) and $(\text{Adam}+\text{L-BFGS})^{2}$ (Optimization Schedule (2)).
Figure 7: L2RE for RBA-PINN, RBF-P, and SAFE-NET on the diffusion PDE using Adam + L-BFGS (left) and $(\text{Adam}+\text{L-BFGS})^{2}$ (right). SAFE-NET, previously ranking second in performance on the diffusion task, is now outperforming the other methods using $(\text{Adam}+\text{L-BFGS})^{2}$.
Figure 8: Average percentage of improvement or deterioration (in L2RE) for each method after switching from Optimization Schedule (1) to Optimization Schedule (2). Figure 12 in the appendix shows percentages for every method on every PDE, demonstrating that SAFE-NET is the only method that produces results in consistent improvement on every PDE after switching to (Adam + L-BFGS)2.
@@ 5.3 Feature Engineering & Spectral Density Analysis
SAFE-NET significantly improves the conditioning of all tested PDEs. We analyze conditioning both at initialization (3k epochs) and after training for a while (100k iterations) using SAFE-NET and PINN with Adam as the optimizer with an initial learning rate of 0.001 and an exponential decay rate of 0.9 every 2k steps (Experimental setup of Figure˜1)
@@
We see that SAFE-NET shifts most eigenvalues values close to 1 at initialization (Figure˜10 top). Post-training spectral density plots (Figure˜10 bottom) reveal dramatic conditioning improvements: the top eigenvalues for the wave problem decrease by a factor of $10^{4}$, while those for the heat and Burgers problems decrease by $10^{2}$ (Figures 25 and 26. Further experiments (Appendix C) show that SAFE-NET reduces both the number and density of large eigenvalues across all problems and all baseline methods (see Table 21)
@@
This improvement in conditioning is a key factor explaining why SAFE-NET features are easier to optimize. Figure 11 compares SAFE-NET and PINNs performance on different loss components for the diffusion and Allen-Cahn PDEs (from the same experiment as Figure 5) as an example. rathore2024challenges argues the residual loss is the main cause of the ill-conditioned loss landscape of PINNs. SAFE-NET significantly improves the conditioning of each loss component, including the residual loss.
@@ 6 Implicit preconditioning in SAFE-NET
Empirical results in Section 5.3 show SAFE-NET improves the conditioning of each problem both at the beginning and end of training. To develop better intuition into this empirical observation for the early stages of training, we consider a simple didactic setting similar to wang2020eigenvector, where the network is given by
@@
$\Theta_{f}(x^{\prime},x)=\nabla_{\theta}f(x^{\prime})^{\top}\nabla f_{\theta}(x).$
@@
where $p$ is the number of trainable parameters and the expectation is taken over the weights at initialization jacot2018neural; liu2020linearity. The neural tangent kernel matrix $\Theta^{\infty}_{f}$ is defined analogously to the tangent kernel matrix.
In the limit $p\rightarrow\infty$, neural net training with $f_{\theta}$ is equivalent to kernel regression with the NTK matrix $\Theta^{\infty}_{f}$. As training is reduced to a kernel regression problem, the convergence speed of gradient-based optimizers is controlled by the conditioning of the NTK matrix $\Theta^{\infty}_{f}$ jacot2018neural; liu2022loss. Thus, a better-conditioned NTK yields a better optimization landscape and faster convergence.
In the context of (11) and under idealized conditions, we continue the following calculations as an example to build intuition into the empirical results suggesting that SAFE-NET features lead to a better conditioned NTK and, so, a better optimization landscape. When $n$ and $p$ are large, the spectrum of $\Theta^{\infty}_{f}$ is closely related to the spectrum of the integral operator $T_{\Theta_{f}}(g)(x)\coloneqq\int_{\Omega}\Theta_{f}(x^{\prime},x)g(x)dx.$ wang2020eigenvector. Thus, we shall obtain control over the $T_{\Theta_{f}}$ spectrum for $f$ given by (11).
@@
$\phi(x,t)=\sum_{i=1}^{N}\sum_{j=1}^{4}c_{j}^{(\ell)}\phi_{j}^{(\ell)}(x,t),$
where $\{\phi_{j}^{(\ell)}\}_{j=1}^{4}$ are defined in equations (5)-(8) and $\ell=1,...,N$ are the number of each type of product features. The kernel is a sum of separable functions as
$\Theta_{u}((x,t),(x^{\prime},t^{\prime}))=\sum_{i=1}^{N}\sum_{j=1}^{4}\big(c_{j}^{(\ell)}\big)^{2}\phi_{j}^{(\ell)}(x,t)\phi_{j}^{(\ell)}(x^{\prime},t^{\prime}).$ (12)
@@
$\int\Theta_{u}((x,t),(x^{\prime},t^{\prime}))\psi(x^{\prime},t^{\prime})\,dx^{\prime}dt^{\prime}=\beta\psi(x,t).$
Substituting $\Theta_{u}((x,t),(x^{\prime},t^{\prime}))$ from equation (12), we get
$\sum_{i=1}^{N}\sum_{j=1}^{4}\big(c_{j}^{(\ell)}\big)^{2}\phi_{j}^{(\ell)}(x,t)\int\phi_{j}^{(\ell)}(x^{\prime},t^{\prime})\psi(x^{\prime},t^{\prime})=\beta\psi(x,t).$ (13)
Set $\alpha_{j}^{(\ell)}:=\int\phi_{j}^{(\ell)}(x^{\prime},t^{\prime})\psi(x^{\prime},t^{\prime})dx^{\prime}dt^{\prime}$ by the domain compactness assumption, so for $\beta\neq 0$, we get
$\psi(x,t)=\frac{1}{\beta}\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(\ell)}\big(c_{j}^{(\ell)}\big)^{2}\phi_{j}^{(\ell)}(x,t).$ (14)
@@
$\frac{1}{\beta}\sum_{i,\ell}^{N}\sum_{j,m}^{4}\big(c_{j}^{(\ell)}\big)^{2}\phi_{j}^{(\ell)}\alpha_{m}^{(\ell)}\big(c_{m}^{(\ell)}\big)^{2}\int\phi_{j}^{(\ell)}(x^{\prime},t^{\prime})\phi_{m}^{(\ell)}(x^{\prime},t^{\prime})$
$=\beta\big(\frac{1}{\beta}\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(\ell)}\big(c_{j}^{(\ell)}\big)^{2}\phi_{j}^{(\ell)}\big)$
Now, $\{\phi_{j}^{(\ell)}\}$ forms an orthonormal basis, as each $\phi$ is chosen from an orthogonal Fourier basis and has unit norm. Thus,
$\frac{1}{\beta}\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(\ell)}\big(c_{j}^{(\ell)}\big)^{4}\phi_{j}^{(\ell)}=\sum_{i=1}^{N}\sum_{j=1}^{4}\alpha_{j}^{(\ell)}\big(c_{j}^{(\ell)}\big)^{2}\phi_{j}^{(\ell)}.$
From this display, we conclude that for each pair $(i,j)$, $\alpha_{j}^{(\ell)}\left(c_{j}^{(\ell)}\right)^{2}=0\quad\text{or}\quad\beta=\left(c_{j}^{(\ell)}\right)^{2}.$ Thus, the eigenvalues of $T_{\Theta_{u}}$ are 0 or equal $(c_{j}^{(\ell)})^{2}$ with corresponding eigenfunction $\psi(x,t)=\alpha_{j}^{(\ell)}\phi_{j}^{(\ell)}(x,t)$. As the non-zero eigenvalues correspond to the directions relevant to learning, we focus on them. Recall the $\{c_{j}^{(\ell)}\}$ are the trainable amplitudes and are set to $1$ at initialization. Hence, at initialization and early stages, we expect the condition number of the matrix $\Theta_{u}$ to be approximately 1. If $u$ were in the infinite width limit, this would imply $\Theta^{\infty}_{u}$ is well-conditioned and fast convergence of gradient-based optimizers. Thus, this idealized example shows the features selected by SAFE-NET can lead to better conditioning and faster convergence. In practice, finite network width and perturbations from domain knowledge features and normalization are expected to introduce some spreading in the eigenvalue distribution. However, experiments show that SAFE-NET maintains a denser eigenvalue distribution around the theoretical prediction from the example (see Figure 10(a)), preserving the core conditioning benefits anticipated by the idealized example and its insights. Early on in training, Figure˜10 shows the eigenvalue distribution is relatively uniform. During training, SAFE-NET gradually adjusts the $c_{j}^{(\ell)}$’s, balancing different frequencies rather than letting any single Fourier mode dominate. As a result, the eigenvalue distribution shifts outwards relatively slowly. Thus, even at the end of the training, Figure˜10 shows the eigenvalue distribution has not changed much from its initialization, so the landscape remains well-conditioned.